OpenStudy (anonymous):

HHHHHEEEEEEEEEELLLLLLLPPPPPPPPPP logsqrt2(4)=x Solve for x

OpenStudy (jhannybean):

something like \(\ 2\log(\sqrt{4}) = x\) ?

OpenStudy (jhannybean):

Or would you mind drawing out your problem? I'm a bit confused with the notation.

OpenStudy (anonymous):

Ya sorry \[\log_\sqrt{2}{4} =x\]

OpenStudy (aum):

Oh, the base is square root 2.

OpenStudy (jhannybean):

For this you ask yourself, to what power of x is \(\ \sqrt{2}\) = 4?

OpenStudy (jhannybean):

 I call this flower power. it essentially means \(\ (\sqrt{2})^x = 4\)